# Factor copula models for non-Gaussian longitudinal data

### v2

@@ 7 Applications
In this section we describe two data sets arose from real world longitudinal studies involving mixed type responses. These data sets are electronically available in the R packages mixAK and lcmm respectively. However in this article, we focus on assessing the temporal dependency of each longitudinal outcomes (continuous/discrete) separately under univariate setup.
@@ 7.1 The PBC 910 data
Figure 1: Subject-specific profiles over time for (i) Serum bilirubin, (ii) Serum albumin and (ii) Hepatomegaly for PBC 910 data set. The dotted lines show average profiles under placebo and D-penicillamine.
@@
Figure 2: Uniform probability plots of the residuals of the best fitting copula models for (i) Serum bilirubin, (ii) Serum albumin and (ii) Hepatomegaly for PBC 910 data set.
@@ 7.2 The PAQUID data
where $\epsilon_{ij}(i,i,d)\sim N(0,1)$. Again, specifications of the factor copula models are similar to Section 6. While fitting these models, we also compare them with the corresponding random effect models. Our aim of this analysis is to describe the decline with age of the global cognitive ability measured by these psychometric tests and to evaluate the association within the longitudinal responses.
@@ 8 Discussion
In this article we introduced factor copula models for arbitary non-Gaussian longitudinal data including covariates. Our proposed models are shown to be effective to model unbalanced longitudinal data with discrete and continuous responses. With parsimonious specification of factor copula parameters our models are easily scalable to accommodate dependence in moderate to high dimensions without computational obstacles. Two stage IFM method is used to estimate the model parameters. Our simulation studies show consistent and reliable estimation of the marginal as well as dependence parameters of the models. We compared our proposed models with some widely used random effect models with similar specification of the fixed covariates.
Factor copula models assumes homogeneous dependence structure for all subjects by construction. But in certain scenarios, this assumption might not be realistic as measurements taken close in time are likely to be more dependent than measurements taken far apart. Therefore, there are further scope of improvements of factor copula models to account for potential time-heterogeneity. For illustrative purposes, we considered two real world data sets which are popular in the joint modeling literature. A quite natural statistical question would be to evaluate the contemporaneous association between each longitudinal responses. This can be addressed in two ways. In the first approach, one can consider another multivariate copula (Gaussian or Student-$t$) to construct joint distribution of the responses at each time point. This approach has been taken by Baghfalaki & Ganjali (2020) with Gaussian copula and a transition model to describe the temporal dependency. Sefidi et al. (2022) used a similar strategy where a bivariate copula is used to associate continuous and ordinal responses while D-vine copulas are used to account for the temporal dependency. Similar to these approaches, one can consider multivariate copulas to associate the subject specific latent variables in the factor copulas, that is correlated latent variables on each factor. Therefore, all dependence and association parameters can be estimated from the second stage of IFM estimation using the joint likelihood. A more elegant approach would be to use correlated random effects in the marginals of each longitudinal responses in order to explain contemporaneous association between each responses. These corresponds to our future investigations in this framework.
Acknowledgments:
Author is grateful to Prof. Sumitra Purkayastha of Applied Statistics Unit, ISI Kolkata and Prof. Sujit K. Ghosh of NC State University for some helpful discussions.

### v3

@@ 7 Applications
In this section, we introduce two datasets derived from real-world longitudinal studies featuring mixed-type responses. These datasets are readily accessible in the R packages mixAK and lcmm, respectively. However, our article concentrates on evaluating the temporal dependency of each longitudinal outcome (whether continuous or discrete) individually within a univariate framework.
@@ 7.1 The PBC 910 data
Figure 1: Subject-specific profiles over time for (i) Serum bilirubin, (ii) Serum albumin and (iii) Hepatomegaly for PBC 910 data set. The dotted lines show average profiles under placebo and D-penicillamine.
@@
Figure 2: Uniform probability plots of the residuals of the best fitting copula models for (i) Serum bilirubin, (ii) Serum albumin and (iii) Hepatomegaly for PBC 910 data set.
@@ 7.2 The PAQUID data
where $\epsilon_{ij}(i.i.d)\sim N(0,1)$. Again, specifications of the factor copula models are similar to Section 6. While fitting these models, we also compare them with the corresponding random effect models. Our aim of this analysis is to describe the decline with age of the global cognitive ability measured by these psychometric tests and to evaluate the association within the longitudinal responses.
@@ 8 Discussion
In this article, we introduced factor copula models tailored for arbitrary non-Gaussian longitudinal data, incorporating covariates. Our proposed models have demonstrated efficacy in modeling unbalanced longitudinal data featuring both discrete and continuous responses. By adopting parsimonious specifications for factor copula parameters, our models exhibit scalability, allowing for seamless accommodation of dependence in moderate to high dimensions without encountering computational hurdles. We employed the two-stage Inference Function for Margins (IFM) method to estimate the model parameters. Through simulation studies, we observed consistent and reliable estimation of both the marginal and dependence parameters of the models. Furthermore, we conducted comparisons between our proposed models and widely used random effect models, employing similar specifications for fixed covariates. There are further scopes of extending this endevour to other bivariate copulas as well.
Factor copula models inherently assume a homogeneous dependence structure for all subjects, which may not always align with reality, especially in scenarios where measurements taken closer in time exhibit stronger dependence compared to those taken further apart. Thus, there is room for further refinement of factor copula models to accommodate potential time-heterogeneity. For illustrative purposes, we examined two widely used real-world datasets prevalent in joint modeling literature. A pertinent statistical inquiry involves evaluating the contemporaneous association between each longitudinal response. This can be approached in two ways. Firstly, one can employ another multivariate copula (such as Gaussian or Student-$t$) to construct the joint distribution of responses at each time point. This methodology has been employed by Baghfalaki & Ganjali (2020), who utilized a Gaussian copula alongside a transition model to describe temporal dependence. Similarly, Sefidi et al. (2022) adopted a similar strategy, employing a bivariate copula to associate continuous and ordinal responses, while leveraging D-vine copulas to account for temporal dependence. Building upon these methodologies, one can incorporate multivariate copulas to associate the subject-specific latent variables in factor copulas, thereby incorporating correlated latent variables on each factor. Consequently, all dependence and association parameters can be estimated through the second stage of IFM estimation using the joint likelihood. Alternatively, a more sophisticated approach involves incorporating correlated random effects in the marginals of each longitudinal response to explain contemporaneous associations between responses. This avenue represents a focal point for our future investigations within this framework.
9 Declarations
Conflict of interest: The author declares no Conflict of interest.